# Spectral-Spatial Transformer with Active Transfer Learning for Hyperspectral Image Classification

### v1

@@ VII Comparison with State-of-the-art Methods
The comparison of HSI classification methods on the SA dataset, as detailed in Table XIII and illustrated in Figure 19, highlights the superior performance of the proposed ATL-SST model across multiple metrics and qualitative benchmarks. Table XIII provides a comprehensive analysis of accuracy, computational efficiency, and robustness under various training, validation, and testing splits, while Figure 19 visually corroborates these findings with ground truth and classification maps.
Quantitatively, ATL-SST demonstrates outstanding accuracy, achieving the highest OA of 99.4864%, AA of 99.7641%, and $\kappa$ of 99.4281%. These metrics significantly surpass the next best-performing models, such as HybViT and SST, which attain OAs of 97.4912% and 97.5318%, respectively. Importantly, ATL-SST maintains this performance consistency across different data splits, achieving perfect accuracy (100%) in multiple scenarios with varying training sample sizes, which underscores its generalization capability. From a computational perspective, ATL-SST achieves a balanced trade-off between accuracy and training efficiency. Its training time of 157.85 seconds is competitive with models like HybViT and SST, while being significantly faster than WaveMamba and PyFormer, which require 1734.50 and 1469.81 seconds, respectively.
@@
Qualitative results in Figure 19 further validate ATL-SST’s efficacy. The classification maps produced by ATL-SST exhibit superior boundary delineation and region specificity, closely mirroring the ground truth. In comparison, models such as SSMamba and WaveMamba show noticeable misclassifications in regions with complex spectral patterns, while transformer-based models like HybViT and PyFormer, although performing well overall, display minor boundary smoothing. ATL-SST, on the other hand, achieves a fine balance between detail preservation and smoothness, resulting in the most accurate visual outputs.
VIII Conclusions and Future Research Directions
@@
This work proposes a novel multi-stage ATL framework that integrates the Spatial-Spectral Transformer (SST) for efficient HSI classification. The proposed framework leverages the strengths of transfer learning, active learning, and spectral-spatial attention mechanisms to address key challenges in HSI classification, such as high spectral dimensionality and limited labeled data. A major contribution of this work is developing an uncertainty-diversity querying mechanism that adaptively selects the most informative and diverse samples for iterative model refinement. This not only optimizes labeling efficiency but also improves the model’s ability to generalize across varying spectral profiles. Additionally, this work introduced a dynamic freezing strategy to selectively freeze and unfreeze SST layers during the transfer learning process, ensuring an optimal balance between computational efficiency and adaptability to new spectral variations. This mechanism significantly reduces computational overhead while preserving critical learned representations, making the approach scalable for real-time or resource-constrained environments. Moreover, the proposed framework incorporates a self-calibration mechanism for spectral and spatial attention layers, which further improves the model’s robustness.
Extensive experimental validation on benchmark HSI datasets demonstrates that the proposed framework consistently outperforms state-of-the-art models, including CNNs, and baseline SSTs, across key metrics such as classification accuracy, robustness, and computational efficiency. The results confirm the effectiveness of integrating ATL with transformer-based models for HSI classification, showcasing its potential for real-world applications where labeling data is expensive and spectral profiles are complex.
@@
While this study makes substantial progress in advancing HSI classification through active transfer learning and spectral-spatial transformers, several avenues for future exploration remain: for instance, future extensions could focus on semi-supervised or self-supervised learning approaches to reduce reliance on labeled data. This would be particularly beneficial in HSI classification, where obtaining labeled samples is costly. Leveraging self-supervised pretext tasks or pseudo-labeling strategies could further enhance the model’s learning capabilities in scenarios with limited labeled data. Although the current uncertainty-diversity-based querying has proven effective, dynamic querying strategies could be explored. For example, combining reinforcement learning with AL may allow the system to learn optimal sampling strategies over time, potentially improving both sample selection and overall learning efficiency. As HSI data can vary significantly between different sensors or geographical regions, exploring domain adaptation techniques within the ATL framework would be a valuable direction. This would allow models to adapt better to unseen datasets or shifts in data distribution, improving the robustness of the classification process in real-world scenarios.

### v2

@@ VII Comparison with State-of-the-art Methods
As shown in Table XI, the results highlight the superior performance of ATL-SST across most experimental settings. For OA, ATL-SST consistently achieves the highest values, with a maximum of 100% in multiple scenarios, underscoring its robustness and effectiveness. Similarly, ATL-SST outperforms all comparative methods in terms of $\kappa$ and AA, achieving top scores of 97.38% and 96.97%, respectively. This reflects the method’s ability to effectively classify diverse classes within the hyperspectral dataset. Notably, ATL-SST demonstrates significant improvement over AGCN and WaveMamba, with AGCN yielding lower performance metrics (e.g., 91.35% OA and 90.64% $\kappa$) in many cases. While HybViT and SST exhibit strong results, particularly in individual settings, their performance is generally less consistent compared to ATL-SST. The training time analysis also reveals ATL-SST’s computational efficiency, with an average training time comparable to SST and WaveFormer, and significantly lower than methods like WaveMamba and PyFormer. The ground truth maps presented in Figure 11 visually depict the classification results for the UH dataset. It is evident from the maps that ATL-SST provides more precise and coherent classification results, with fewer misclassified pixels and better-defined class boundaries compared to the comparative methods. For instance, methods like AGCN and WaveMamba exhibit noticeable misclassifications and noise, whereas ATL-SST maintains high spatial and spectral accuracy.
The results presented in Table XII and Figure 12 comprehensively compare the performance. In terms of accuracy, ATL-SST consistently outperforms all other methods, achieving the highest OA of 98.8778%, AA of 98.1747%, and $\kappa$ of 98.5130%. These results are a significant improvement over the next best-performing methods, such as WaveMamba, which achieves an OA of 96.1941%. From a computational perspective, ATL-SST maintains competitive efficiency, with a training time of 124.78s, comparable to HybViT and SST while significantly outperforming computationally intensive models like WaveMamba and WaveFormer, which require 1382.70s and 112.68s, respectively. Qualitative results presented in Figure 12 further validate the quantitative findings. The classification maps generated by ATL-SST are visually superior, exhibiting precise boundary delineation and accurate region classifications that closely match the ground truth. In contrast, competing methods, such as AGCN and SSMamba, show noticeable errors, particularly in regions with complex class boundaries or high spectral variability.
@@
The comparison of HSI classification methods on the SA dataset, as detailed in Table XIII and illustrated in Figure 13, highlights the superior performance of the proposed ATL-SST model. Quantitatively, ATL-SST demonstrates outstanding accuracy, achieving the highest OA of 99.4864%, AA of 99.7641%, and $\kappa$ of 99.4281%. These metrics significantly surpass the next best-performing models, such as HybViT and SST, which attain OAs of 97.4912% and 97.5318%, respectively. From a computational perspective, ATL-SST achieves a balanced trade-off between accuracy and training efficiency. Its training time of 157.85s is competitive with models like HybViT and SST, while being significantly faster than WaveMamba and PyFormer, which require 1734.50s and 1469.81s, respectively. Qualitative results in Figure 13 further validate ATL-SST’s efficacy. The classification maps produced by ATL-SST exhibit superior boundary delineation and region specificity, closely mirroring the ground truth. In comparison, models such as SSMamba and WaveMamba show noticeable misclassifications in regions with complex spectral patterns, while transformer-based models like HybViT and PyFormer, although performing well overall, display minor boundary smoothing.
VIII Ablation Study
To evaluate the individual contributions of each module within the SST-ATL framework, we conduct a comprehensive ablation study on three representative datasets: PU, UH, and SA. The ablation analysis is divided into two categories: component-level evaluation and query strategy assessment.
Component-Level Ablation
First, to measure the effect of AL, we replace the hybrid uncertainty-diversity sampling strategy with random sampling. Second, to isolate the impact of diversity sampling, we retain only uncertainty-based selection. Third, to test the influence of selective model adaptation, we disable dynamic freezing and allow all transformer layers to be uniformly fine-tuned. Fourth, we examine the value of the self-calibrated attention mechanism by reverting to standard MHSA without uncertainty-guided refinement.
The results presented in Table XIV indicate that each component significantly contributes to the final performance of the model. The full SST-ATL achieves the highest OA, AA, and $\kappa$ across all datasets. The removal of AL causes the most severe degradation, OA drops by 2.66%, 2.65%, and 2.70% for PU, UH, and SA respectively, highlighting the effectiveness of the proposed hybrid sampling. Excluding diversity sampling also leads to performance drops, particularly in the AA metric, reflecting reduced class coverage. The absence of dynamic freezing results in slightly reduced accuracy but increased training instability, while the removal of self-calibrated attention degrades spatial-spectral feature alignment, as evidenced by reduced $\kappa$ values.
TABLE XIV: Component-level ablation results.
Variant PU UH SA
OA AA $\kappa$ OA AA $\kappa$ OA AA $\kappa$
Full SST-ATL 98.88 98.17 98.51 97.58 96.97 97.38 99.49 99.76 99.43
w/o Active Learning 96.22 94.60 95.00 94.93 93.48 94.51 96.79 97.90 96.90
w/o Diversity 97.14 96.13 96.47 95.72 94.36 95.08 97.85 98.40 97.86
w/o Dynamic Freezing 97.31 96.23 96.60 96.14 94.72 95.52 98.12 98.83 98.10
w/o Self-Calibrated Attention 97.53 96.74 97.12 96.80 95.93 96.61 98.71 99.12 98.63
Query Strategy Evaluation
In addition to component analysis, we evaluate the effectiveness of our hybrid uncertainty-diversity sampling strategy against four widely used alternatives: random sampling, entropy-based uncertainty sampling, margin-based uncertainty sampling, and diversity-only sampling. All strategies are tested across four AL rounds (250 to 1000 labeled samples) on the PU, UH, and SA datasets.
As shown in Table XV, our proposed hybrid query strategy outperforms all others in every setting. The hybrid method achieves up to 98.14% OA on PU, 97.29% on UH, and 98.97% on SA with only 1000 labeled samples. The advantage is particularly evident in the early rounds, where the model learns from as few as 250 samples. For example, at 250 labels, the hybrid method exceeds random sampling by over 3% OA on all datasets, showcasing its superior label efficiency and informativeness. These results emphasize that incorporating both uncertainty and spectral-spatial diversity into the query process leads to faster convergence, better generalization, and significantly lower annotation costs than existing AL approaches.
TABLE XV: Comparison of query strategies across AL iterations.
Strategy PU (OA%) UH (OA%) SA (OA%)
250 500 750 1000 250 500 750 1000 250 500 750 1000
Random Sampling 90.81 93.54 95.02 96.22 88.13 91.95 93.41 94.93 91.74 94.27 95.89 96.79
Entropy-Based Uncertainty 91.54 94.38 95.73 97.14 89.97 93.22 94.78 95.72 92.90 95.83 96.61 97.85
Margin Sampling 91.23 94.07 95.49 96.80 89.51 92.68 94.33 95.31 92.46 95.30 96.20 97.43
Diversity-Only Sampling 92.14 94.96 96.34 97.04 90.62 93.75 95.11 95.96 93.33 96.08 96.95 98.00
Proposed (Hybrid) 93.92 96.23 97.51 98.14 92.01 95.04 96.30 97.29 94.71 97.42 98.13 98.97
IX Conclusions and Future Research Directions
@@
This work proposes a novel multi-stage ATL framework that integrates the SST for efficient HSI classification. The proposed framework leverages the strengths of transfer learning, AL, and SST to address key challenges in HSI classification, such as high spectral dimensionality and limited labeled data. A major contribution of this work is developing an uncertainty-diversity querying mechanism that adaptively selects the most informative and diverse samples for iterative model refinement. This not only optimizes labeling efficiency but also improves the model’s ability to generalize across varying spectral profiles. Additionally, this work introduced a dynamic freezing strategy to selectively freeze and unfreeze SST layers during the transfer learning process, ensuring an optimal balance between computational efficiency and adaptability to new spectral variations. This mechanism significantly reduces computational overhead while preserving critical learned representations, making the approach scalable for resource-constrained environments.
@@
While this study makes substantial progress in advancing HSI classification through ATL and SST, several avenues for future exploration remain: for instance, future extensions could focus on self-supervised learning approaches to reduce reliance on labeled data. This would be particularly beneficial in HSI classification, where obtaining labeled samples is costly. Leveraging self-supervised pretext tasks or pseudo-labeling strategies could further enhance the model’s learning capabilities in scenarios with limited labeled data. Although the current uncertainty-diversity-based querying has proven effective, dynamic querying strategies could be explored. For example, combining reinforcement learning with AL may allow the system to learn optimal sampling strategies over time, potentially improving both sample selection and overall learning efficiency. As HSI data can vary significantly between different sensors or geographical regions, exploring domain adaptation techniques within the ATL framework would be a valuable direction. This would allow models to adapt better to unseen datasets or shifts in data distribution, improving the robustness of the classification process in real-world scenarios.